if 600ml of a juice drink is 15% orange juice, how much orange juice should be added to make a drink that is 25% orange juice
Answer by mananth(16949)

--------- percent ---------------- quantity
orange drink 15----------------600
orange juice 100----------------x
Total 25 ---------------- 600+x
....
15*600+100*x=25(600+x)
9000+100 x=15000 +25x
100x-25x =15000-9000
75x=6000 /7
/75
x=80 ml orange juice